# Massive Activations in Large Language Models

### v1

@@ Abstract
We observe an empirical phenomenon in Large Language Models (LLMs)—very few activations exhibit significantly larger values than others (e.g., 100,000 times larger). We call them massive activations. First, we demonstrate the widespread existence of massive activations across various LLMs and characterize their locations. Second, we find their values largely stay constant regardless of the input, and they function as indispensable bias terms in LLMs. Third, these massive activations lead to the concentration of attention probabilities to their corresponding tokens, and further, implicit bias terms in the self-attention output. Last, we also study massive activations in Vision Transformers. [^1] Code is available at https://github.com/locuslab/massive-activations.
@@ 2.2 Which Feature and Sequence Dimensions?
Mixtral-8x7B. For this particular model, massive activations lie in two feature dimensions, i.e., 2070 and 3398. For sequence dimensions, we find that they are associated with the starting token, delimiter tokens and also certain word tokens, e.g., token “and” and token “of”. Figure 3 showcases these patterns within Mixtral-8x7B. Generally, for inputs of 4096 tokens in length, these tokens are predominantly located in the early part of sequence.
@@ 4 Effects on Attention
In this section, we study the internal mechanism of massive activations in LLMs, particularly in relation to self-attention.
@@ 5 Massive Activations in Vision Transformers
[Uncaptioned image] Figure 12: Three largest activation magnitudes and the median magnitude at each layer in CLIP and DINOv2 ViT-L.
@@
[Uncaptioned image] Figure 14: Average attention of the [CLS] token.
@@ 6 Related Work
Biases in self-attention. There can be various notion of biases in the self-attention mechanism. First, simple additive bias terms can be used in linear layers for computing the query, key and value states (Namazifar et al., 2023). Second, position biases can be inserted in self-attention to encode positional information of each token (Su et al., 2021, Press et al., 2021). There are also variants of biases with manually designed softmax operator (Evan, 2023, Bondarenko et al., 2023). Our work reveals that LLMs, even with standard self-attention, would impose implicit bias components in the attention computation through massive activations.
7 Conclusion
@@ A.1 Pretrained LLMs
In Section 2, we have demonstrated massive activations in LLaMA2-7B, LLaMA2-13B and Mixtral-8x7B. In this section, we evaluate more pretrained LLMs which cover a wide range of model families. We illustrate massive activations in LLaMA2-70B, Phi-2, Mistral-7B (Jiang et al., 2023), MPT-7B (MosaicML, 2023) and Falcon-7B (Almazrouei et al., 2023). The results are presented in Figure 15, 16, 17, 18 and 19.
@@ A.2 Fine-tuned LLMs
Our results so far are focused on pretrained LLMs. However, a significant application of LLMs lies in their use for chat purposes. Instruction fine-tuning (Ouyang et al., 2022) is essential for developing models capable of generating coherent responses to questions. In this part, we demonstrate massive activations in these fine-tuned models. We evaluate fine-tuned models from models in LLaMA2 and Mistral. The results are shown in Figure 20, 21, 22 and 23.
@@ A.3 BOS Token <s>
In Figure 2425 and 26, we show massive activations in LLaMA2-7B, LLaMA2-13B and Mixtral-8x7B, with the same input sequences as in Section 2. We find that massive activations persist with a prepended BOS token. In LLaMA2-7B and LLaMA2-13B, the locations of massive activations, i.e., sequence and feature dimensions, are not altered. However, for Mixtral-8x7B, some massive activations shift to the BOS token <s>. We leave the study on how the BOS token <s> affects the positions of massive activations for future work.
@@ B.4 Explicit Attention Biases
    The first one is a special case of our proposed formulation in Equation 3, where both $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ are zero vectors. Equation 6 shows the computation of this variant of self-attention. This is also equivalent to the previous proposed Softmax-off-by-one (Evan, 2023).
@@ C.3 Layer-Level Analysis
Figure 40, 41 and 42 detail the layer-level analysis results for all ViTs examined in this paper (also summarized in Table 8). Different from LLMs, some ViTs do not exhibit massive activations, e.g., MAE ViT-B/L and DINOv2 ViT-S. For ViTs where we observe massive activations, e.g., CLIP ViT-L and DINOv2 ViT-L, the trend across layers differs from LLMs. For instance, in the case of DINOv2 ViT-L, massive activations are observed in the later stages of this model but are absent in the output of the final layer.

### v2

@@ Abstract
We observe an empirical phenomenon in Large Language Models (LLMs)—very few activations exhibit significantly larger values than others (e.g., 100,000 times larger). We call them massive activations. First, we demonstrate the widespread existence of massive activations across various LLMs and characterize their locations. Second, we find their values largely stay constant regardless of the input, and they function as indispensable bias terms in LLMs. Third, these massive activations lead to the concentration of attention probabilities to their corresponding tokens, and further, implicit bias terms in the self-attention output. Last, we also study massive activations in Vision Transformers. Code is available at https://github.com/locuslab/massive-activations. [^1] Published as a conference paper in First Conference on Language Modeling (COLM), 2024.
@@ 2.2 Which Feature and Sequence Dimensions?
Mixtral-8x7B. For this model, massive activations lie in two feature dimensions, i.e., 2070 and 3398. For sequence dimensions, we find that they are associated with the starting token, delimiter tokens and also certain word tokens, e.g., token “and” and token “of”. These word tokens tend to be conjunctions and prepositions, representing relatively few semantics. Figure 3 showcases these patterns in Mixtral-8x7B. Generally, for inputs of 4096 tokens in length, these tokens are predominantly located in the early part of sequence.
@@ 4 Effects on Attention
In this section, we explore and study the internal mechanism of massive activations in LLMs, particularly in relation to self-attention.
@@ 5 Massive Activations in Vision Transformers
Figure 12: Three largest activation magnitudes and the median magnitude at each layer in CLIP and DINOv2 ViT-L.
CLIP ViT-L, layer 13
@@
Original 75.5
Set to zero 59.8
Set to mean 75.5
Table 4: Intervention analysis of massive activations in CLIP ViT-L.
@@
Figure 14: Average attention of the [CLS] token.
DINOv2-reg with 4 registers
@@
Original 81.9 84.8 86.3 87.0
Fix-Reg-Mean 81.7 85.0 86.2 87.0
Table 5: We fix all register features at every layer to their means and evaluate the intervened ViTs.
@@ 6 Related Work
Biases in self-attention. There can be various notion of biases in the self-attention mechanism. First, simple additive bias terms can be used in linear layers for computing the query, key and value states (Namazifar et al., 2023). Second, position biases can be inserted in self-attention to encode positional information of each token (Su et al., 2021, Press et al., 2021). There are also variants of biases with manually designed softmax operator (Miller, 2023, Bondarenko et al., 2023, Hu et al., 2024). Our work reveals that LLMs, even with standard self-attention formulation, would impose implicit bias components in the attention computation through massive activations.
7 Conclusion and Discussion
@@
We discuss some practical implications and future directions of this work. First, the presence of activations with large magnitudes has been widely known as a major challenge in effectively quantizing LLMs (Dettmers et al., 2022, Xiao et al., 2023a). This paper identifies a new type of outlier activations in LLMs, and we hope our findings will be of value to research on LLM compression. Second, attention maps that allocate excessive attention probabilities to a few fixed tokens may be undesirable for mechanistic interpretability (Olsson et al., 2022). Our proposed attention formulation could make the resulting attention maps in LLMs more interpretable, and potentially benefit downstream applications (Darcet et al., 2023). Finally, our investigation of the new attention formulation is focused on its effects on massive activations, and our experiments were limited to a small GPT-2 model due to computational resource constraints. It would be interesting to see how our results generalize to models at larger scales, and how our attention formulation could affect the training stability (Wortsman et al., 2023) of modern LLMs.
@@ A.1 Pretrained LLMs
In Section 2, we have demonstrated massive activations in LLaMA2-7B, LLaMA2-13B and Mixtral-8x7B. In this section, we evaluate more pretrained LLMs which cover a wide range of model families. We illustrate massive activations in LLaMA2-70B, LLaMA3 (Dubey et al., 2024), Phi-2, Mistral-7B (Jiang et al., 2023), MPT-7B (MosaicML, 2023) and Falcon-7B (Almazrouei et al., 2023). The results are presented in Figure 15, 16, 17, 18, 19, 20 and 21.
@@
Figure 16: Massive activations in LLaMA3-8B.
Figure 17: Massive activations in LLaMA3-70B.
@@ A.2 Fine-tuned LLMs
Our results so far are focused on pretrained LLMs. However, a significant application of LLMs lies in their use for chat purposes. Instruction fine-tuning (Ouyang et al., 2022) is essential for developing models capable of generating coherent responses to questions. In this part, we demonstrate massive activations in these fine-tuned models. We evaluate fine-tuned models from models in LLaMA2 and Mistral. The results are shown in Figure 22, 23, 24 and 25.
@@ A.3 BOS Token <s>
In Figure 2627 and 28, we show massive activations in LLaMA2-7B, LLaMA2-13B and Mixtral-8x7B, with the same input sequences as in Section 2. We find that massive activations persist with a prepended BOS token. In LLaMA2-7B and LLaMA2-13B, the locations of massive activations, i.e., sequence and feature dimensions, are not altered. However, for Mixtral-8x7B, some massive activations shift to the BOS token <s>. We leave the study on how the BOS token <s> affects the positions of massive activations for future work.
@@ B.4 Explicit Attention Biases
Figure 37: Attention distribution in default GPT-2 and GPT-2 with explicit attention bias.
In Figure 37, we visualize the attention distribution in both default GPT-2 and GPT-2 with explicit attention biases, where we plot the average attention probability over 50 sentences each with 30 tokens. First, we find that our observations on the relationship between massive activations and attention concentration hold for the default GPT-2 model. Second, for the GPT-2 model with explicit attention bias, most of the attention probability is assigned to the extra $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ vectors we inserted. Intriguingly, this also holds for initial layers as well (e.g., layer 1), suggesting the strong need for LLMs to form this attention concentration pattern during pretraining.
@@
    The first one is a special case of our proposed formulation in Equation 3, where both $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ are zero vectors. Equation 6 shows the computation of this variant of self-attention. This is also equivalent to the previous proposed Softmax-off-by-one (Miller, 2023).
@@ C.3 Layer-Level Analysis
Figure 43, 44 and 45 detail the layer-level analysis results for all ViTs examined in this paper (also summarized in Table 8). Different from LLMs, some ViTs do not exhibit massive activations, e.g., MAE ViT-B/L and DINOv2 ViT-S. For ViTs where we observe massive activations, e.g., CLIP ViT-L and DINOv2 ViT-L, the trend across layers differs from LLMs. For instance, in the case of DINOv2 ViT-L, massive activations are observed in the later stages of this model but are absent in the output of the final layer.
@@ Appendix D Models and Datasets
LLaMA3 LLaMA3-8B 32 4096 32 meta-llama/Meta-Llama-3-8B
LLaMA3-70B 80 8192 64 meta-llama/Meta-Llama-3-70B